# Visual analysis of bivariate dependence between continuous random variables

### v1

@@ Introduction
The main contributions of this paper are: (1) explaining, through copula function theory, the inadequacies of scatter plots for dependence visualization, and the preference for measures like Spearman’s rank correlation over Pearson’s correlation; (2) advocating for rank plots as more suitable tools for analyzing statistical dependence; (3) providing interpretation guidelines for common patterns in rank plots; and (4) introducing a novel graphical ensemble that we call a dplot, for performing a comprehensive analysis of the relationship between two continuous random variables.
The remainder of this paper is structured as follows: A review of relevant literature is presented in Section 1, followed by a discussion on how marginal distributions influence scatter plots under an identical dependence relationship in Section 2. In Section 3 we briefly delve into copula function theory, while in 4 we revisit key concepts of dependence, association measures, and graphical representations. In Section 5 we introduce the concept of a dplot, and in Section 6 we show their utility through several examples employing real data. Finally, Section 7 contains the main conclusions and a discussion.
@@ 3 Bivariate copulas
Basic results and properties about copula functions may be found in [28]. For example, the underlying copula of a random vector is invariant under continuous strictly increasing transformations of its random variables:
@@ 4.1 Quadrant dependence
Another significant aspect concerning copulas is the ability to ascertain the proximity to PQD, NQD and independence. As we have seen, a bivariate uniform distribution of observations from $\left(F_{X}(X),F_{Y}(Y)\right)$ indicates independence. Therefore, a departure from uniformity indicates some type of dependence. The maximum deviation from independence can be established by applying Sklar’s theorem (1) to the Fréchet-Hoeffding bounds [21, 12] for bivariate joint distributions:
@@ 4.3.2 Gluing copulas
$C_{1,2,\theta}(u,v)=\begin{cases}\,\theta C_{1}\left(\frac{u}{\theta},v\right),\text{ if }0\leq u\leq\theta,\\ \,(1-\theta)C_{2}\left(\frac{u-\theta}{1-\theta},v\right),\text{ if }\theta\leq u\leq 1.\end{cases}$ (11)
@@ Acknowledgements
This work was supported by a grant from Universidad Rey Juan Carlos (2023/SOLCON-132212).

### v2

@@ Introduction
The main contributions of this paper are: (1) explaining, through copula function theory, the inadequacies of scatter plots for dependence visualization, and the preference for measures like Spearman’s rank correlation over Pearson’s correlation; (2) advocating for rank plots as more suitable tools for analyzing statistical dependence; (3) providing interpretation guidelines for common patterns in rank plots; (4) introducing a novel graphical ensemble that we call a dplot, for performing a comprehensive analysis of the relationship between two continuous random variables; and (5) all the Julia programming code and data sets used for calculations and generating figures is available for reproducibility.
The remainder of this paper is structured as follows: A review of relevant literature is presented in Section 1, followed by a discussion on how marginal distributions influence scatter plots under an identical dependence relationship in Section 2. In Section 3 we briefly delve into copula function theory, while in 4 we revisit key concepts of dependence, association measures, and graphical representations. In Section 5 we introduce the concept of a dplot, and in Section 6 we show their utility through several examples employing real data. Finally, Section 7 contains the main conclusions and a discussion, and Section 8 provides a link to the Julia programming code for reproducibility of all calculations and figures.
@@ 3 Bivariate copulas
Basic results and properties about copula functions may be found in [29] or [24]. For example, the underlying copula of a random vector is invariant under continuous strictly increasing transformations of its random variables:
@@ 4.1 Quadrant dependence
Another significant aspect concerning copulas is the ability to ascertain the proximity to PQD, NQD and independence. As we have seen, a uniform distribution on $[0,1]^{2}$ of observations from $\left(F_{X}(X),F_{Y}(Y)\right)$ indicates independence. Therefore, a departure from uniformity indicates some type of dependence. The maximum deviation from independence can be established by applying Sklar’s theorem (1) to the Fréchet-Hoeffding bounds [21, 12] for bivariate joint distributions:
@@ 4.3.2 Gluing copulas
$C_{1,2,\theta}(u,v)=\begin{cases}\,\theta C_{1}\left(\frac{u}{\theta},v\right),\text{ if }0\leq u\leq\theta,\\ \,(1-\theta)C_{2}\left(\frac{u-\theta}{1-\theta},v\right)+\theta v,\text{ if }\theta\leq u\leq 1.\end{cases}$ (11)
@@ 7 Conclusions and discussion
8 Reproducibility
All the Julia programming code and data sets used for calculations and generating figures is available for reproducibility at https://github.com/aerdely/visualdep
@@ Acknowledgements
This work was partially supported by a grant from Universidad Nacional Autónoma de México (PAPIIT IN104425) and by a grant from Universidad Rey Juan Carlos (2023/SOLCON-132212).